OpenStudy (anonymous):

could you tell me if this is correct then? problem=> (x+3)^2 answer is => (x+3)(x+3) = xx+x*3+3*x+3*3 = x^4+12 ?

OpenStudy (anas.p):

no, your first step was correct xx+x*3+3*x+3*3 but xx=x^2\[xx=x^{2}\] x*3 and 3*x both will give the ans as 3x each. Together they are 3x+3x=6x. 3*3=9 So,\[x^{2}+6x=9\]

OpenStudy (anonymous):

gee thanks i think i have that one undrstood then... appreciate your help

OpenStudy (anas.p):

cool